# HV-BEV: Decoupling Horizontal and Vertical Feature Sampling for Multi-View 3D Object Detection

### v2

@@ IV-D Ablation Study
Main Ablations In this subsection, we focus on validating the effectiveness of the proposed enhancement strategies. As shown in Tab. IV, the vanilla BEVFormer achieves an mAP of 37.0% and an NDS of 47.9%. Incorporating the Vertical Adaptive Height-Aware (VHA) sampling strategy improves these metrics by 0.5% and 0.3%, respectively. Adding the Horizontal Cross-view Feature Aggregation (HCA) module further boosts mAP and NDS by 2.2% and 1.9%, respectively. When both strategies are applied simultaneously, the mAP and NDS reach 39.9% and 50.2%, representing gains of 2.9% and 2.3%, respectively. Additionally, significant improvements are observed across other metrics as well.
TABLE IV: Main ablation results of HV-BEV on nuScenes val set. ”VHA” and ”HCA” denotes for vertical adaptive height-aware sampling and horizontal cross-view feature aggregation, respectively.
@@
[…]
Ablations on VHA In the VHA module, the number of height bins determines the sampling density of 3D reference points. Setting too few height bins (e.g., 4) degrade the model into uniform height sampling, similar to the vanilla BEVFormer. Conversely, an excessive number of height bins may cause reference points to overly concentrate within a narrow height range. Therefore, selecting an optimal number of height bins is crucial. Tab. V summarizes the experimental results for different height bin configurations. When the number of height bins is set to 8, the model achieves the best performance. The performance with 6 bins is only slightly lower than with 8 bins, but increasing the number of bins beyond 8 leads to a notable decline in performance. We infer that this decline occurs because the supervision for the height distribution prediction is based on the 3D center of the object. As the number of bins increases and becomes denser, the 3D reference points are sampled closer to the object center, resulting in higher probabilities being assigned to bins near the center. This concentration on the central region may force the model to overlook the broader height range, thereby degrading overall performance.
TABLE V: Effectiveness of the number of height bins in VHA module on nuScenes val set.
@@
Ablations on HCA In the HCA module, each 3D reference point constructs a horizontal dynamic graph centered on itself, with each node in the graph aggregating relevant features from its respective location. Theoretically, increasing the number of graph nodes allows the aggregation of richer features but also incurs higher computational costs. Determining the optimal number of neighboring nodes to balance accuracy and computational efficiency is a critical consideration. Tab. VI presents the model’s performance and memory usage under various settings for the number of neighbors. Interestingly, performance degrades when the number of neighbors becomes excessively large, potentially due to over-smoothing. We adopt 4 neighbors as the default configuration, which not only reduces memory usage but also maintains competitive accuracy.
TABLE VI: Effectiveness of the number of neighbors in HCA module on nuScenes val set.
@@
2 0.386 0.489 $\sim 11700$M
4 0.399 0.502 $\sim 13900$M
8 0.402 0.504 $\sim 17150$M
16 0.379 0.480 $\sim 22950$M
Effectiveness of decoupled sampling method This study enhances feature aggregation through height-aware reference point sampling along the vertical axis and graph-based cross-view perception in the horizontal plane. However, it is generally acknowledged that increasing the number of reference points for a single query, without considering computational cost, inherently allows the sampling of more image features, thereby improving detection accuracy. The purpose of this section is to validate the effectiveness of the proposed feature sampling strategies, rather than attributing performance improvements solely to the increased number of reference points. Tab. VII reports the results of experiments conducted on the baseline model under the $small$ configuration. These tests examine the effect of increasing the number of uniformly sampled 3D reference points per query ($N_{ref}=4,8,16$) on detection accuracy. The results indicate that increasing $N_{ref}$ to 8 yields only minor performance improvements, while further increasing it to 16 results in a decline across all metrics. We attribute this decline to the potential redundancy or irrelevance introduced by simply adding more uniformly sampled reference points, which may weaken the impact of critical features and disrupt the model’s learning process. These findings further validate the effectiveness of the proposed decoupled horizontal and vertical feature sampling strategies.
TABLE VII: Effectiveness of increasing the number of 3D reference points within each query on nuScenes val set.

### v3

@@ IV-E Ablation Study
Main Ablations In this subsection, we focus on validating the effectiveness of the proposed enhancement strategies. As shown in Tab. VI, the vanilla BEVFormer achieves an mAP of 37.0% and an NDS of 47.9%. Incorporating the Vertical Adaptive Height-Aware (VHA) sampling strategy improves these metrics by 0.5% and 0.3%, respectively. Adding the Dynamic Horizontal Cross-view Feature Aggregation (DHCA) module further boosts mAP and NDS by 2.2% and 1.9%, respectively. When both strategies are applied simultaneously, the mAP and NDS reach 39.9% and 50.2%, representing gains of 2.9% and 2.3%, respectively. Additionally, significant improvements are observed across other metrics as well.
TABLE VI: Main ablation results of HV-BEV on nuScenes val set. ”VHA” and ”DHCA” denotes for vertical adaptive height-aware sampling and dynamic horizontal cross-view feature aggregation, respectively.
@@
      0.370       0.479       0.721       0.279       0.407       0.436       0.220
      0.375       0.482       0.716       0.280       0.405       0.431       0.217
      0.392       0.498       0.709       0.279       0.399       0.409       0.208
      0.399       0.502       0.706       0.279       0.395       0.400       0.203
Ablations on VHA In the VHA module, the number of height bins determines the sampling density of 3D reference points. Setting too few height bins (e.g., 4) degrade the model into uniform height sampling, similar to the vanilla BEVFormer. Conversely, an excessive number of height bins may cause reference points to overly concentrate within a narrow height range. Therefore, selecting an optimal number of height bins is crucial. Tab. VII summarizes the experimental results for different height bin configurations. When the number of height bins is set to 8, the model achieves the best performance. The performance with 6 bins is only slightly lower than with 8 bins, but increasing the number of bins beyond 8 leads to a notable decline in performance. We infer that this decline occurs because the supervision for the height distribution prediction is based on the 3D center of the object. As the number of bins increases and becomes denser, the 3D reference points are sampled closer to the object center, resulting in higher probabilities being assigned to bins near the center. This concentration on the central region may force the model to overlook the broader height range, thereby degrading overall performance. Although this issue can be alleviated by increasing the number of $N_{ref}$, doing so inevitably incurs higher computational overhead, as each 3D reference point and its neighboring points require feature sampling from the image. Moreover, overly dense sampling may lead to performance degradation due to over-smoothing.
TABLE VII: Effectiveness of the number of height bins in VHA module on nuScenes val set.
@@
Ablations on DHCA In the DHCA module, each 3D reference point constructs a horizontal dynamic graph centered on itself, with each node in the graph aggregating relevant features from its respective location. Theoretically, increasing the number of neighboring nodes allows the aggregation of richer features but also incurs higher computational costs. Determining the optimal number of neighboring nodes to balance accuracy and computational efficiency is a critical consideration. Tab. VIII presents a comparison of the model’s accuracy, memory consumption, and inference speed under different settings of the number of neighboring nodes. Interestingly, performance degrades when the number of neighbors becomes excessively large, potentially due to over-smoothing. The inference speed remains comparable when the number of neighbors is set to 2 or 4; however, a significant drop in FPS is observed when the number exceeds 8. We adopt 4 neighbors as the default configuration, which not only reduces memory usage but also maintains competitive accuracy.
TABLE VIII: Effectiveness of the number of neighbors in DHCA module on nuScenes val set.
@@
2 0.386 0.489 $\sim 11700$M 9.87
4 0.399 0.502 $\sim 13900$M 9.08
8 0.402 0.504 $\sim 17150$M 7.25
16 0.379 0.480 $\sim 22950$M 3.04
Effectiveness of decoupled sampling method This study enhances feature aggregation through height-aware reference point sampling along the vertical axis and graph-based cross-view perception in the horizontal plane. However, it is generally acknowledged that increasing the number of reference points for a single query, without considering computational cost, inherently allows the sampling of more image features, thereby improving detection accuracy. The purpose of this section is to validate the effectiveness of the proposed feature sampling strategies, rather than attributing performance improvements solely to the increased number of reference points. Tab. IX reports the results of experiments conducted on the baseline model under the $small$ configuration. These tests examine the effect of increasing the number of uniformly sampled 3D reference points per query ($N_{ref}=4,8,16$) on detection accuracy. The results indicate that increasing $N_{ref}$ to 8 yields only minor performance improvements, while further increasing it to 16 results in a decline across all metrics. We attribute this decline to the potential redundancy or irrelevance introduced by simply adding more uniformly sampled reference points, which may weaken the impact of critical features and disrupt the model’s learning process. These findings further validate the effectiveness of the proposed decoupled horizontal and vertical feature sampling strategies.
TABLE IX: Effectiveness of increasing the number of 3D reference points within each query on nuScenes val set.